OpenStudy (anonymous):

Please could I get some help with this last question I have been waiting for 2 hours please! I just need it in (P/X)/DP/DX form please! Demand function 128 ------ +2 x^2 Quality demand x = 4 Find the price elasticity of demand for the demand function at the indicated x-value.

OpenStudy (anonymous):

@JackleBee

OpenStudy (anonymous):

you mean you want to get rid of the fraction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

shoot, did you try logs?

OpenStudy (anonymous):

nope, not sure what that is.

OpenStudy (anonymous):

y doesn't equal to anything at the moment right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Ive tried different methods, but not getting a clear answer

OpenStudy (anonymous):

which methods did you try? you might have missed a step

OpenStudy (anonymous):

I tried 2x= 128+ 2 to get rid of the fraction, but I got the wrong answer

OpenStudy (anonymous):

lol why 2x?

OpenStudy (anonymous):

I just tried to move the x^2 over to one side then solve 128/2 to get 64

OpenStudy (anonymous):

128 ------ +2 = y x^2 did you try square root?

OpenStudy (anonymous):

like move it, and then square root?

OpenStudy (anonymous):

man, i'm stumped on this one, I'll get this other guy who might help. ;p @satellite73

OpenStudy (anonymous):

sorry dude, but I want to see how this is solved as well xD

OpenStudy (anonymous):

same here. I know the process, but the fraction is throwing me off

OpenStudy (anonymous):

@satellite73 you think you can help us sir? and yeah, that fraction xD

OpenStudy (anonymous):

I think I remember @satellite73

OpenStudy (anonymous):

I'm trying to get this problem into (P/X)/DP/DX form, but the fraction is throwing me off

jimthompson5910 (jim_thompson5910):

What is P when x = 4 ?

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

err 10

jimthompson5910 (jim_thompson5910):

so you'll have P/x = 10/4 = 2.5

jimthompson5910 (jim_thompson5910):

dP/dx is simply the derivative of P with respect to x

jimthompson5910 (jim_thompson5910):

think of it as P = 128x^(-2) + 2 and use the power rule

OpenStudy (anonymous):

\[\frac{ -256 }{ x^3 }\]

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now I think you plug in x = 4 into dP/dx

OpenStudy (anonymous):

I get -4

OpenStudy (anonymous):

I got 2.5/-4 = -5/8

jimthompson5910 (jim_thompson5910):

so, (P/x)/(dP/dx) = (2.5)/(-4) = -0.625 essentially the price elasticity of demand is -62.5% for every increase of the price, the demand drops by 62.5%. I think this is only true near the value x = 4

OpenStudy (anonymous):

Wow thank you so much ! :) I have been stuck in that problem for the longest.

jimthompson5910 (jim_thompson5910):

np